OpenStudy (anonymous):

what is 4x+6y=-2 and 3x-y=9 of elimination

OpenStudy (anonymous):

do we have to find the value of x and y??

OpenStudy (anonymous):

what occurs when the second row is multiplied by 6 and the two equations are added?

OpenStudy (anonymous):

4x+6y=-2 ---(1) 3x-y=9 ---(2) simplify (1), 2x+3y = -1 ---(3) (2)x3: 9x -3y =27 ---(4) <--- this step is used to produce 3y which (3) has it too (3)+(4) : 11x = 26 <-- eliminate 3y in (3) and (4) x = 26/11 Put this into (3), 2(26/11) +3 y = -1 y= -21/11

OpenStudy (anonymous):

then 2nd row will become 18x-6y=54 and after adding two equations we will get value of x =26/11

OpenStudy (anonymous):

thanks you guys .